# Unconditional verification of quantum computation with classical light

### v1

@@ III Verification of quantum computation with coherent light
    it is less than $m_{0}^{(j)}+m_{1}^{(j)}$ for all $1\leq j\leq N$.
We derive an upper bound on the acceptance probability $p_{\rm acc}$ in each case one by one. By taking the maximum of these two upper bounds, we can derive an upper bound on $p_{\rm acc}$ of our physically-classical verification protocol.
We first consider the first case (i). Let $j^{\ast}$ such that $\tilde{m}_{0}^{(j^{\ast})}\geq m_{0}^{(j^{\ast})}+m_{1}^{(j^{\ast})}$. From the Hoeffding inequality H63 , the probability of the prover not being rejected in step (d) of the $j^{\ast}$-th repetition of Protocol 2 is
${\rm Pr}\left[\tilde{m}_{0}^{(j^{\ast})}\leq me^{-\alpha^{2}}\left(1+\cfrac{\alpha^{2}}{2}\right)\right]$ (14)
$\leq$ ${\rm Pr}\left[m\cfrac{1+\alpha^{2}}{e^{\alpha^{2}}}-\left(m_{0}^{(j^{\ast})}+m_{1}^{(j^{\ast})}\right)\geq me^{-\alpha^{2}}\cfrac{\alpha^{2}}{2}\right]$
$\leq$ ${\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right),$ (15)
where we have used $\tilde{m}_{0}^{(j^{\ast})}\geq m_{0}^{(j^{\ast})}+m_{1}^{(j^{\ast})}$ and the fact that $m_{0}^{(j^{\ast})}+m_{1}^{(j^{\ast})}$ converges to $me^{-\alpha^{2}}(1+\alpha^{2})$ to derive the first and second inequalities, respectively FNM2 . Since $\tilde{m}_{0}^{(j^{\ast})}\leq me^{-\alpha^{2}}(1+\alpha^{2}/2)$ is a necessary condition to be accepted, from Eq. (15),
$p_{\rm acc}\leq{\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right).$ (16)
In the second case (ii), $\tilde{m}_{0}^{(j)}<m_{0}^{(j)}+m_{1}^{(j)}$ holds, and hence the input of the I1DC protocol must include at least a single state whose actual photon number is one or zero. Note that we can assume that the malicious prover is not rejected in step (d) of Protocol 2 because our purpose is to derive an upper bound on $p_{\rm acc}$.
We first consider the case where a single state whose actual photon number is one is included in the input of the I1DC protocol. Let it be the $l^{\ast}$-th input state whose polarization angle is $\sigma_{l^{\ast}}^{(j)}\in\{0,\pi/2,\pi,3\pi/2\}$. In general, the malicious prover applies any CPTP map to the $(m-\tilde{m}_{0}^{(j)})$ input state that are used for the I1DC protocol in the case of the honest prover. We interpret it as that the $l^{\ast}$-th input state is converted to some quantum state by using the other $(m-\tilde{m}_{0}^{(j)}-1)$ input state as ancillary states. This interpretation implies that, immediately before step (e) in Protocol 2, the prover’s state is
@@
with some CPTP map $\mathcal{E}$. Since the verifier chooses each polarization angle independently in step (a) of Protocol 2, and the actual photon number of the $l^{\ast}$-th input state is one, $\mathcal{E}$ does not depend on $\sigma_{l^{\ast}}^{(j)}$ but can depend on the other input states’ polarization angles $\{\sigma_{l}^{(j)}\}_{1\leq l\leq m-\tilde{m}_{0}^{(j)},l\neq l^{\ast}}$. To mimic the honest prover, in steps (e) and (f) of Protocol 2, the malicious prover generates and sends the measurement outcomes $\vec{o}^{\ (j)}\equiv\{o_{l^{\prime}}^{(j)}\}_{l^{\prime}=1}^{m-\tilde{m}_{0}^{(j)}-1}\in\{0,1\}^{m-\tilde{m}_{0}^{(j)}-1}$ by applying some additional CPTP map to the quantum state in Eq. (17). For simplicity, we define $|+_{\theta}\rangle\equiv(|0\rangle+e^{i\theta}|1\rangle)/\sqrt{2}$ for any real number $\theta$. The malicious prover’s state is finally
@@
where $p_{\sigma_{l^{\ast}}^{(j)}}(\vec{o}^{\ (j)})$ is the probability of outputting $\vec{o}^{\ (j)}$, and $\mathcal{E}_{\vec{o}^{\ (j)}}$ is a quantum operation to be applied when the measurement outcomes are $\vec{o}^{\ (j)}$. The subscript of $p_{\sigma_{l^{\ast}}^{(j)}}(\vec{o}^{\ (j)})$ just represents that the probability, in general, depends on $|+_{\sigma_{l^{\ast}}^{(j)}}\rangle$ and does not mean that the prover’s CPTP map is constructed by using the value of $\sigma_{l^{\ast}}^{(j)}$. In fact, there exits a CPTP map $\mathcal{F}_{1}^{(j)}$ such that
@@
for any $\sigma_{l^{\ast}}^{(j)}$. This is due to the fact that the verifier keeps the value of $\sigma_{l^{\ast}}^{(j)}$ private, and the actual photon number of the $l^{\ast}$-th input state is one, i.e., the information of $\sigma_{l^{\ast}}^{(j)}$ is only contained in $|+_{\sigma_{l^{\ast}}^{(j)}}\rangle$. For any $\vec{o}^{\ (j)}$ and $\{\sigma_{l}^{(j)}\}_{l\neq l^{\ast}}$, there exist $\sigma\in\{0,\pi/2,\pi,3\pi/2\}$ and $c\in\{1,-1\}$ such that
@@
We next consider the case where a vacuum state is included in the input of the I1DC protocol. Let it be originated from the $k^{\ast}$-th coherent state whose polarization angle is $\theta_{k^{\ast}}^{(j)}$. Since the verifier selects each polarization angle independently, the malicious prover’s state $\rho^{(j)}$ at the end of Protocol 2 does not depend on $\theta_{k^{\ast}}^{(j)}$ and hence $\varphi^{(j)}$. Remember that $\varphi^{(j)}$ is just a variable replacement of $\theta_{k^{\ast}}^{(j)}$. By using the CPTP map $\mathcal{F}_{2}^{(j)}$ that replaces any quantum state with the fixed state $\rho^{(j)}$, the prover’s final state is
@@
By combining the above arguments, when the deviation of the prover is independent in each repetition, the malicious prover’s state after the $N$-th repetition in Protocol 2 is
@@
where $\mathcal{F}^{(j)}\in\{\mathcal{F}_{1}^{(j)},\mathcal{F}_{2}^{(j)}\}$ for all $1\leq j\leq N$. As shown in Appendix B, the similar argument holds even when the prover’s attack is collective, i.e., the prover simultaneously handles all photons in all repetitions to deceive the verifier. Therefore, $p_{\rm acc}\leq 1-b/2$. From this upper bound and Eq. (16), when the prover is malicious,
$p_{\rm acc}\leq{\rm max}\left\{{\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right),1-\cfrac{b}{2}\right\}.$ (27)
@@
$-{\rm max}\left\{{\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right),1-\cfrac{b}{2}\right\}$
$\geq$ $1-\cfrac{a}{2}-\cfrac{1}{2Nf(|x|)}-\left(1-\cfrac{b}{2}\right)$
@@
which is the inverse of a polynomial in $|x|$. Thus, our physically-classical verification protocol efficiently distinguishes the honest and malicious provers.
@@ IV Conclusion & discussion
We have proposed an efficient verification protocol that removes the necessity of the communication of qubits. Since all apparatuses required for the verifier are a telecom-band laser with linear optical elements and a classical computer, our results would facilitate the realization of verifiable quantum computers.
To further improve the practicality of our protocol, we discuss the phase randomization implemented for the coherent state. In several papers (e.g., Ref. CZLM15 ), the difficulty of the continuous phase randomization was pointed out. By calculating the fidelity between the continuous-phase-randomized (see Eq. (4)) and discrete-phase-randomized coherent states, we evaluate how many classical bits would be required for the phase randomization. For simplicity, let
@@
Let $p_{\rm acc}$ and $q_{\rm acc}$ be the acceptance probabilities of our physically-classical verification protocol with continuous-phase-randomized and discrete-phase-randomized coherent states, respectively. Eq. (29) implies that $|p_{\rm acc}-q_{\rm acc}|\leq(N-1)/[4Nf(|x|)]$ is sufficient to that our protocol correctly distinguishes the honest and malicious provers. From Ref. HQ12 that studies the fidelity and trace distance for any quantum states in an infinite-dimensional separable complex Hilbert space, $|p_{\rm acc}-q_{\rm acc}|\leq\sqrt{1-F_{\rm min}}$, and hence using $R\geq e^{2}+1$ results in $R=\lceil\log{[32mN^{3}f(|x|)^{2}/(N-1)^{2}]}+1\rceil$. Since $m=\lceil 2e^{2}\log{(2N^{2}f(|x|))}\rceil$, this calculation shows that a logarithmic of logarithmic number of classical bits is sufficient for the phase randomization.
@@
The RBSP protocol DKL12 was improved or modified in Refs. ZL16 ; ZL18 ; NHS19 ; PYZMDZLW22 . As another direction to improve our protocol, it would be interesting to consider their applicability to the verification of quantum computation. Furthermore, as with the implementation security ZNC24 of QKD, it would be important to devise the verification protocols under several imperfections such as the channel loss and noises and the correlation between coherent states. An efficient way KLMO24 that removes trusted quantum state preparations and measurements from verification protocols may be useful for this purpose.
@@
In this paper, we propose a physically-classical verification protocol by combining the protocols in Refs. M20 and DKL12 . On the other hand, there are compilers that make blind quantum computing protocols verifiable M18 ; KKLMO22 . Blind quantum computation is a secure protocol such that a user can delegate universal quantum computation to a remote quantum computer without disclosing the user’s input, quantum algorithm, and output. Although Ref. M18 seems to implicitly assume the perfect blindness (i.e., the perfect security), the blind quantum computing protocol with coherent states proposed in Ref. DKL12 does not satisfy the perfect blindness. This is why it would not be directly applied to Ref. DKL12 to obtain a physically-classical verification protocol. When we apply Ref. KKLMO22 to Ref. DKL12 , the resultant protocol should require a polynomial number of communication rounds between the verifier and prover, while our protocol is one round because $N$ repetitions in Protocol 2 can be done in parallel, and the prover can send $\{n_{k}^{(j)}\}_{k,j}$ and $\{w,z\}_{w,z}$ simultaneously. Our protocol successfully reveals an advantage of the transmission of coherent light in the sense that if the ultimate protocol in Fig. 1 with a constant number of rounds can be constructed, then ${\sf BQP}$ is contained in the third level of the polynomial hierarchy FHM18 . This set containment is considered to be unlikely from an oracle separation between ${\sf BQP}$ and ${\sf PH}$ RT22 .
@@
It would be interesting to devise physically-classical verification protocols by modifying other existing verification protocols (e.g., Ref. B18 ) and then combining it with the RBSP protocol DKL12 .
@@ ACKNOWLEDGMENTS
In this appendix, we derive the malicious prover’s quantum state after the $N$-th repetition in Protocol 2 under any collective attack. Since we consider the second case (ii) where the number $\tilde{m}_{0}^{(j)}$ of vacuum states is less than $m_{0}^{(j)}+m_{1}^{(j)}$ for all $1\leq j\leq N$, the I1DC protocol in the $j$-th repetition has a state whose actual photon number is zero or one as an input. Let it be the $l^{\ast}_{j}$-th input state whose polarization angle is $\sigma_{l^{\ast}_{j}}^{(j)}$. For simplicity, we define $\vec{\sigma}\in\{0,\pi/2,\pi,3\pi/2\}^{(m-1)N}$ as the string of all the polarization angles except for $\{\sigma_{l^{\ast}_{j}}^{(j)}\}_{j=1}^{N}$. In general, the prover’s final state can be written as
@@
where $p_{\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)},\vec{\sigma}}(\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)})$ is the probability of outputting $\{\vec{o}^{\ (j)}\}_{j=1}^{N}$, and $\mathcal{E}_{\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)},\vec{\sigma}}$ is a quantum operation to be applied when the measurement outcomes are $\{\vec{o}^{\ (j)}\}_{j=1}^{N}$. The subscript $\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)}$ of $p_{\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)},\vec{\sigma}}(\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)})$ just represents that the probability can depend on each $|+_{\sigma_{l^{\ast}_{j}}^{(j)}}\rangle$ if the $l^{\ast}_{j}$-th input qubit of the I1DC protocol in the $j$-th repetition is not a vacuum state, and it does not mean that the prover’s CPTP map is constructed by using the values of $\{\sigma_{l^{\ast}_{j}}^{(j)}\}_{j=1}^{N}$. Another subscript $\vec{\sigma}$ of $p_{\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)},\vec{\sigma}}(\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)})$ and $\mathcal{E}_{\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)},\vec{\sigma}}$ means that they may depend on $\vec{\sigma}$ because the corresponding coherent states may include more than one photon. Since the verifier keeps the values of $\{\sigma_{l^{\ast}_{j}}^{(j)}\}_{j=1}^{N}$ private, and the actual photon number of each $|+_{\sigma_{l^{\ast}_{j}}^{(j)}}\rangle$ is one or zero, as with Eq. (19), there exists a CPTP map $\mathcal{F}$ such that
@@
for any $\{\sigma_{l^{\ast}_{j}}^{(j)}\}_{j=1}^{N}$. Remember that when the actual photon number of $|+_{\sigma_{l^{\ast}_{j}}^{(j)}}\rangle$ is zero, the probability $p_{\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)},\vec{\sigma}}(\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)})$ does not depend on $\sigma_{l^{\ast}_{j}}^{(j)}$.
@@
Since this quantum state can be prepared at the end of step 1 in Protocol 1, the inequality $p_{\rm acc}\leq 1-b/2$ holds.

### v2

@@ III Verification of quantum computation with coherent light
    $m_{0}^{(j)}+m_{1}^{(j)}$ is more than the threshold value for all $1\leq j\leq N$.
Let $p_{\rm(i)}$ and $p_{\rm(ii)}$ denote the probabilities of the first and second cases occurring, respectively. We also define $p_{{\rm acc}|{\rm(i)}}$ and $p_{{\rm acc}|{\rm(ii)}}$ as probabilities of the prover being accepted in the first and second cases, respectively. By using these notations, the acceptance probability $p_{\rm acc}$ of our physically classical verification protocol is
$p_{\rm acc}$ $=$ $p_{{\rm acc}|{\rm(i)}}p_{\rm(i)}+p_{{\rm acc}|{\rm(ii)}}p_{\rm(ii)}\leq p_{\rm(i)}+p_{{\rm acc}|{\rm(ii)}}.$ (14)
We derive an upper bound on $p_{\rm acc}$ by evaluating $p_{\rm(i)}$ and $p_{{\rm acc}|{\rm(ii)}}$ one by one.
We first evaluate $p_{\rm(i)}$. From the Hoeffding inequality H63 , for all $j$,
${\rm Pr}\left[m_{0}^{(j)}+m_{1}^{(j)}>me^{-\alpha^{2}}\left(1+\cfrac{\alpha^{2}}{2}\right)\right]$ (15)
$=$ $1-{\rm Pr}\left[m_{0}^{(j)}+m_{1}^{(j)}\leq me^{-\alpha^{2}}\left(1+\cfrac{\alpha^{2}}{2}\right)\right]$
$=$ $1-{\rm Pr}\left[m\cfrac{1+\alpha^{2}}{e^{\alpha^{2}}}-\left(m_{0}^{(j)}+m_{1}^{(j)}\right)\geq me^{-\alpha^{2}}\cfrac{\alpha^{2}}{2}\right]$ (16)
$\geq$ $1-{\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right),$ (17)
where we have used the fact that $m_{0}^{(j)}+m_{1}^{(j)}$ converges to $me^{-\alpha^{2}}(1+\alpha^{2})$ to derive the last inequality FNM2 . Therefore,
$p_{\rm(i)}$$\leq$ $1-\left[1-{\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right)\right]^{N}$ (18)
$\leq$ $N{\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right).$ (19)
We next evaluate $p_{{\rm acc}|{\rm(ii)}}$. Let $\tilde{m}_{0}^{(j)}$ be the number of vacuum states calculated from the measurement outcomes sent by the malicious prover in step (c) of Protocol 2. The inequality $\tilde{m}_{0}^{(j)}<m_{0}^{(j)}+m_{1}^{(j)}$ has to hold for all $j$ to maximize $p_{{\rm acc}|{\rm(ii)}}$ because the malicious prover is definitely rejected if $\tilde{m}_{0}^{(j)}\geq m_{0}^{(j)}+m_{1}^{(j)}>me^{-\alpha^{2}}(1+\alpha^{2}/2)$. This implies that the input of the I1DC protocol must include at least a single state whose actual photon number is one or zero.
We first consider the case where a single state whose actual photon number is one is included in the input of the I1DC protocol. Let it be the $l^{\ast}$th input state whose polarization angle is $\sigma_{l^{\ast}}^{(j)}\in\{0,\pi/2,\pi,3\pi/2\}$. In general, the malicious prover applies any CPTP map to the $(m-\tilde{m}_{0}^{(j)})$ input states, which are used for the I1DC protocol in the case of an honest prover. We interpret it as that the $l^{\ast}$th input state is converted to some quantum state by using the other $(m-\tilde{m}_{0}^{(j)}-1)$ input states as ancillary states. This interpretation implies that, immediately before step (e) in Protocol 2, the prover’s state is
@@
with some CPTP map $\mathcal{E}$. Since the verifier chooses each polarization angle independently in step (a) of Protocol 2, and the actual photon number of the $l^{\ast}$th input state is one, $\mathcal{E}$ does not depend on $\sigma_{l^{\ast}}^{(j)}$ but can depend on the other input states’ polarization angles $\{\sigma_{l}^{(j)}\}_{1\leq l\leq m-\tilde{m}_{0}^{(j)},l\neq l^{\ast}}$. To mimic the honest prover, in steps (e) and (f) of Protocol 2, the malicious prover generates and sends the measurement outcomes $\vec{o}^{\ (j)}\equiv\{o_{l^{\prime}}^{(j)}\}_{l^{\prime}=1}^{m-\tilde{m}_{0}^{(j)}-1}\in\{0,1\}^{m-\tilde{m}_{0}^{(j)}-1}$ by applying some additional CPTP map to the quantum state in Eq. (20). For simplicity, we define $|+_{\theta}\rangle\equiv(|0\rangle+e^{i\theta}|1\rangle)/\sqrt{2}$ for any real number $\theta$. The malicious prover’s state is finally
@@
where $p_{\sigma_{l^{\ast}}^{(j)}}(\vec{o}^{\ (j)})$ is the probability of outputting $\vec{o}^{\ (j)}$, and $\mathcal{E}_{\vec{o}^{\ (j)}}$ is a quantum operation to be applied when the measurement outcomes are $\vec{o}^{\ (j)}$. The subscript of $p_{\sigma_{l^{\ast}}^{(j)}}(\vec{o}^{\ (j)})$ just represents that the probability, in general, depends on $|+_{\sigma_{l^{\ast}}^{(j)}}\rangle$; it does not mean that the prover’s CPTP map is constructed by using the value of $\sigma_{l^{\ast}}^{(j)}$. In fact, there exists a CPTP map $\mathcal{F}_{1}^{(j)}$ such that
@@
for any $\sigma_{l^{\ast}}^{(j)}$. This is because the verifier keeps the value of $\sigma_{l^{\ast}}^{(j)}$ private and the actual photon number of the $l^{\ast}$th input state is one, i.e., the information of $\sigma_{l^{\ast}}^{(j)}$ is only contained in $|+_{\sigma_{l^{\ast}}^{(j)}}\rangle$. For any $\vec{o}^{\ (j)}$ and $\{\sigma_{l}^{(j)}\}_{l\neq l^{\ast}}$, there exist $\sigma\in\{0,\pi/2,\pi,3\pi/2\}$ and $c\in\{1,-1\}$ such that
@@
We next consider the case where a vacuum state is included in the input of the I1DC protocol. Let the vacuum state be originated from the $k^{\ast}$th coherent state whose polarization angle is $\theta_{k^{\ast}}^{(j)}$. Since the verifier selects each polarization angle independently, the malicious prover’s state $\rho^{(j)}$ at the end of Protocol 2 does not depend on $\theta_{k^{\ast}}^{(j)}$ and hence $\varphi^{(j)}$. Remember that $\varphi^{(j)}$ is just a variable replacement of $\theta_{k^{\ast}}^{(j)}$. By using the CPTP map $\mathcal{F}_{2}^{(j)}$ that replaces any quantum state with the fixed state $\rho^{(j)}$, the prover’s final state is
@@
By combining the above arguments, when the deviation of the prover is independent in each repetition, the malicious prover’s state after the $N$th repetition in Protocol 2 is
@@
where $\mathcal{F}^{(j)}\in\{\mathcal{F}_{1}^{(j)},\mathcal{F}_{2}^{(j)}\}$ for all $1\leq j\leq N$. As shown in Appendix B, a similar argument holds even when the prover’s attack is collective, i.e., the prover simultaneously handles all photons in all repetitions to deceive the verifier. Therefore, $p_{{\rm acc}|{\rm(ii)}}\leq 1-b/2$. From this upper bound and Eqs. (14) and (19), when the prover is malicious,
$p_{\rm acc}\leq 1-\cfrac{b}{2}+N{\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right).$ (30)
@@
$-\left[1-\cfrac{b}{2}+N{\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right)\right]$
$\geq$ $\cfrac{b-a}{2}-\cfrac{1}{2Nf(|x|)}$
@@
which is the inverse of a polynomial in $|x|$. Thus, our physically classical verification protocol efficiently distinguishes the honest and malicious provers.
@@ IV Conclusion & discussion
We have proposed an efficient verification protocol that removes the necessity of qubit communication. Since all apparatuses required for the verifier are a telecom-band laser with linear optical elements and a classical computer, our results would facilitate the realization of verifiable quantum computers.
To further improve the practicality of our protocol, we discuss the phase randomization implemented for the coherent state. Several papers (e.g., Ref. CZLM15 ) have pointed out the difficulty of the continuous phase randomization. By calculating the fidelity between the continuous-phase-randomized (see Eq. (4)) and discrete-phase-randomized coherent states, we evaluate how many classical bits would be required for the phase randomization. For simplicity, let
@@
Let $p_{\rm acc}$ and $q_{\rm acc}$ be the acceptance probabilities of our physically classical verification protocol with continuous-phase-randomized and discrete-phase-randomized coherent states, respectively. Eq. (32) implies that $|p_{\rm acc}-q_{\rm acc}|\leq(N-1)/[4Nf(|x|)]$ is sufficient for our protocol to correctly distinguish between honest and malicious provers. From Ref. HQ12 that studies the fidelity and trace distance for any quantum states in an infinite-dimensional separable complex Hilbert space, $|p_{\rm acc}-q_{\rm acc}|\leq\sqrt{1-F_{\rm min}}$, and hence using $R\geq e^{2}+1$ results in $R=\lceil\log{[32mN^{3}f(|x|)^{2}/(N-1)^{2}]}+1\rceil$. Since $m=\lceil 2e^{2}\log{(4N^{2}f(|x|))}\rceil$, this calculation shows that a log of log number of classical bits is sufficient for the phase randomization.
@@
The RBSP protocol DKL12 was improved or modified in Refs. ZL16 ; ZL18 ; NHS19 ; PYZMDZLW22 . As another direction to improve our protocol, it would be interesting to consider their applicability to the verification of quantum computation. Furthermore, as with the implementation security ZNC24 of QKD, it would be important to devise the verification protocols under several imperfections such as channel loss and noises and the correlation between coherent states. An efficient way KLMO24 that removes trusted quantum state preparations and measurements from verification protocols may be useful for this purpose.
@@
In this paper, we have proposed a physically classical verification protocol by combining the protocols in Refs. M20 and DKL12 . On the other hand, there are compilers that make blind quantum computing protocols verifiable M18 ; KKLMO22 . Blind quantum computation is a secure protocol such that a user can delegate universal quantum computation to a remote quantum computer without disclosing the user’s input, quantum algorithm, and output. Although Ref. M18 seems to implicitly assume perfect blindness (i.e., perfect security), the blind quantum computing protocol with coherent states proposed in Ref. DKL12 does not have perfect blindness. This is why it could not be directly applied to Ref. DKL12 to obtain a physically classical verification protocol. When we apply Ref. KKLMO22 to Ref. DKL12 , the resultant protocol should require a polynomial number of communication rounds between the verifier and prover, while our protocol is one round because $N$ repetitions in Protocol 2 can be done in parallel, and the prover can send $\{n_{k}^{(j)}\}_{k,j}$ and $\{w,z\}_{w,z}$ simultaneously. Our protocol successfully reveals an advantage of the transmission of coherent light in the sense that if the ultimate protocol in Fig. 1 with a constant number of rounds can be constructed, then ${\sf BQP}$ is contained in the third level of the polynomial hierarchy FHM18 . This set containment is considered to be unlikely due to an oracle separation between ${\sf BQP}$ and ${\sf PH}$ RT22 .
@@
It would be interesting to devise physically classical verification protocols by modifying other existing verification protocols (e.g., Ref. B18 ) and then combining them with the RBSP protocol DKL12 .
@@ ACKNOWLEDGMENTS
In this appendix, we derive the malicious prover’s quantum state after the $N$th repetition in Protocol 2 under any collective attack. Since we consider case (ii), we can assume that the number $\tilde{m}_{0}^{(j)}$ of vacuum states is less than $m_{0}^{(j)}+m_{1}^{(j)}$ for all $1\leq j\leq N$. Therefore, the I1DC protocol in the $j$th repetition has a state whose actual photon number is zero or one as an input. Let it be the $l^{\ast}_{j}$th input state whose polarization angle is $\sigma_{l^{\ast}_{j}}^{(j)}$. For simplicity, we define $\vec{\sigma}\in\{0,\pi/2,\pi,3\pi/2\}^{(m-1)N}$ as the string of all the polarization angles except for $\{\sigma_{l^{\ast}_{j}}^{(j)}\}_{j=1}^{N}$. In general, the prover’s final state can be written as
@@
where $p_{\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)},\vec{\sigma}}(\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)})$ is the probability of outputting $\{\vec{o}^{\ (j)}\}_{j=1}^{N}$, and $\mathcal{E}_{\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)},\vec{\sigma}}$ is a quantum operation to be applied when the measurement outcomes are $\{\vec{o}^{\ (j)}\}_{j=1}^{N}$. The subscript $\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)}$ of $p_{\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)},\vec{\sigma}}(\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)})$ just represents that the probability can depend on each $|+_{\sigma_{l^{\ast}_{j}}^{(j)}}\rangle$ if the $l^{\ast}_{j}$th input qubit of the I1DC protocol in the $j$th repetition is not a vacuum state; it does not mean that the prover’s CPTP map is constructed by using the values of $\{\sigma_{l^{\ast}_{j}}^{(j)}\}_{j=1}^{N}$. The other subscript, $\vec{\sigma}$, of $p_{\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)},\vec{\sigma}}(\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)})$ and $\mathcal{E}_{\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)},\vec{\sigma}}$ means that they may depend on $\vec{\sigma}$ because the corresponding coherent states may include more than one photon. Since the verifier keeps the values of $\{\sigma_{l^{\ast}_{j}}^{(j)}\}_{j=1}^{N}$ private and the actual photon number of each $|+_{\sigma_{l^{\ast}_{j}}^{(j)}}\rangle$ is one or zero, as with Eq. (22), there exists a CPTP map $\mathcal{F}$ such that
@@
for any $\{\sigma_{l^{\ast}_{j}}^{(j)}\}_{j=1}^{N}$. Recall that when the actual photon number of $|+_{\sigma_{l^{\ast}_{j}}^{(j)}}\rangle$ is zero, the probability $p_{\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)},\vec{\sigma}}(\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)})$ does not depend on $\sigma_{l^{\ast}_{j}}^{(j)}$.
@@
Since this quantum state can be prepared at the end of step 1 in Protocol 1, the inequality $p_{{\rm acc}|{\rm(ii)}}\leq 1-b/2$ holds.